Grade 9Most Scoring Topics in IIT JEE

p and Q are two elements which forms P2 Q3 and PQ2 it 0.15 mole of P2Q3 weight 5.9 grams and 0.15 mole of PQ2 9.3gm. what are atomic weights of PandQ

SANDEEP
8 Years agoGrade 9

1 Answer

Arun
8 Years ago

0.15 moles of P2Q3 weighs 15.9 g

Therefore , 1 moles of P2Q3 weigh 106 g

Now, 0.15 moles of PQ2 weigh 9.3 gn

1 mole of PQ2 will weigh 62 g

Now, 2P + 3Q = 106 --- (i)

Also, P + 2Q = 62----(ii)

2P + 4Q = 124 --- (iii)

(iii) - (i)

2P + 4Q - ( 2P + 3 Q) = 124 - 106

Q = 18 , P = 26

Ans: Q = 18 , P = 26